OpenStudy (anonymous):

how do i solve: 49.51cosx - 30sinx = 1.22625

OpenStudy (anonymous):

\[49\times51timescosx \div \sqrt{\left( 49\times51 \right)^{2}+30^{2}}+30timessinx \div \sqrt{\left( 49\times51 \right)^{2}+30^{2}}=1.22625\div \sqrt{\left( 49\times51 \right)^{2}+30^{2}}\]

OpenStudy (anonymous):

\[\cosx =\left(\begin{matrix}30 \\ \sqrt{\left( 49\times51 \right)^{2}+30^{2}}\end{matrix}\right)\] \[sinx =\left(\begin{matrix}49\times51 \\ \sqrt{\left( 49\times51 \right)^{2}+30^{2}}\end{matrix}\right)\]

OpenStudy (anonymous):

i'm not following any of that.... 49x51 is not in the question